# Uniform Norm Error Estimates for 2D Turning Point Problem

### v1

@@ 4 Estimates for convection and reaction terms involving discrete Green’s function
$\leq C\|IF_{12}-F_{12}\|_{L^{\infty}(\Omega_{c})}\varepsilon^{-1/2}\;\|G_{mn}\|_{1,\varepsilon}$
$\leq CN\varepsilon^{-1/2}[h_{x}^{2}\|F_{12,xx}\|_{L^{\infty}(\Omega_{c})}+h_{x}h_{y}\|F_{12,xy}\|_{L^{\infty}(\Omega_{c})}$
$\quad+h_{y}^{2}\|F_{12,yy}\|_{L^{\infty}(\Omega_{c})}]\leq\frac{CN}{\varepsilon^{1/2}}\left[\frac{\varepsilon^{\alpha}}{\varepsilon^{2}N^{2}}+\frac{\varepsilon^{\alpha}}{\varepsilon^{3/2}N^{2}}+\frac{\varepsilon^{\alpha}}{\varepsilon N^{2}}\right]$
$\leq C\left[\frac{\varepsilon^{\alpha-\frac{5}{2}}}{N}\right]\leq\frac{C}{N}\quad\text{when}\;\alpha>5/2.$ (4.25)
Hence, when source point $(x_{m},y_{n})\in\Omega_{c}$ then we have
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{c}}\leq\frac{C}{N}.$
Similar arguments we can show that when source point $(x_{m},y_{n})\in\Omega_{f,x}$, then we have
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{c}}\leq\frac{C}{N}.$
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{c}}\leq\frac{C}{N}.$
$(Iu-u,b_{1}G_{mn,x})_{\Omega}\leq C\frac{\log^{2}(1/\varepsilon^{3/2})}{N}.$
@@
Similarly arguments yields that when source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, then we have
$(Iu-u,b_{1,x}G_{mn})_{\Omega}\leq C\frac{\log^{2}(1/\varepsilon^{3/2})}{N}.$
@@
Lemma 5.5 If the source point $(x_{m},y_{n})$ lies in $\Omega_{c}\cup\Omega_{f,x}$, then $(c(Iu-u),G_{mn}).\leq C\frac{\log^{2}(1/\varepsilon^{3/2})}{N}.$
[…]
$(c(Iu-u),G_{mn})_{\Omega}$ $=(c(Iu-u),G_{mn})_{\Omega_{c}}$
$+(c(Iu-u),G_{mn})_{\Omega_{f,x}}$
$+(c(Iu-u),G_{mn})_{\Omega_{f,y}}$
$+(c(Iu-u),G_{mn})_{\Omega_{f,xy}}$ (4.26)
Similar arguments yields that when source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$ then the reaction term bounds by
$(c(Iu-u),G_{mn})_{\Omega}\leq C\frac{\log^{2}(1/\varepsilon^{3/2})}{N}.$
Theorem 4
@@
When $\alpha\geq 5/2$ and $\beta\geq 2/3$ and $u_{h}$ denotes the approximate solution of the problem, and the source point $(x_{m},y_{n})$ only lies in the coarse region $\Omega_{c}$ and $x$-layer region $\Omega_{f,x}$. Building upon the preceding analysis, the following estimate holds
$|(u-u_{h})(x_{m},y_{n})|\leq C\frac{(\log(1/\varepsilon^{3/2}))^{3}}{N}.$
Proof. At the beginning, we aim to consider $|(Iu-u_{h})(x_{m},y_{n})|$
Let $(x_{m},y_{n})$ be the mesh point then
$(Iu-u_{h})(x_{m},y_{n})=(\delta_{mn},Iu-u_{h})$ (4.27)
@@
Where $\delta_{mn}$ is the Kronecker delta function defined on mesh point $(x_{m},y_{n})$. Let $G_{mn}$ be the discrete Green’s function corresponding to the differential operator coming from problem (1.2) and the discrete Dirac-Delta function $\delta_{mn}$, then
$(Iu-u_{h})(x_{m},y_{n})=(\delta_{mn},Iu-u_{h})=B(Iu-u_{h},G_{mn})$ (4.28)
where
$B(u,v)=\varepsilon\int_{\Omega}\nabla u.\nabla v\;dx+\int_{\Omega}b_{1}u_{x}v\;dx+\int_{\Omega}cuv\;dx$
$B(u,v)=\varepsilon(\nabla u,\nabla v)+(b_{1}u_{x},v)+(cu,v)$ (4.29)
$\implies(Iu-u_{h})(x_{m},y_{n})=B(Iu-u,G_{mn})+B(u-u_{h},G_{mn})$ (4.30)
But by the Galerkin’s orthogonality, we can write that
$B(u-u_{h},G_{mn})=0$ (4.31)
$\implies(Iu-u_{h})(x_{m},y_{n})=B(Iu-u,G_{mn})$ (4.32)
$\implies(Iu-u_{h})(x_{m},y_{n})=\varepsilon\int_{\Omega}\nabla(Iu-u).\nabla G_{mn}\;dxdy\\ +\int_{\Omega}b_{1}(Iu-u)_{x}G_{mn}\;dxdy+\int_{\Omega}c(Iu-u)G_{mn}\;dxdy$
$\implies(Iu-u_{h})(x_{m},y_{n})=\varepsilon(\nabla(Iu-u),\nabla G_{mn})+(b_{1}(Iu-u)_{x},G_{mn})\\ +(c(Iu-u),G_{mn})$ (4.33)
In this theorem, we aim to bound the bilinear form B(.,.) appearing in (4.29) with its arguments the interpolation error uIuu-Iu and the discrete Green’s function $G_{mn}.$ We decompose the solution $u$ into regular and singular parts, where we note that the singular part denotes the part of the solution that contains layers. Therefore, the singular part can be decomposed further into three more regions, namely $F_{1}$, $F_{2}$, and $F_{12}$, where $F_{1}$ contains the interior layer, $F_{2}$ contains the boundary layer, and $F_{12}$ contains both the interior and boundary layers in the corner region.
To establish error estimates, we have to estimate the term on the R.H.S. of (4.33), we begin in the first term of the R.H.S. of (4.33) is $\varepsilon(\nabla(Iu-u),\nabla G_{mn})$ and decompose the solution $u=R+F_{1}+F_{2}+F_{12}$, we obtain
$\varepsilon(\nabla(u-Iu),\nabla G_{mn})=\varepsilon(\nabla((R+F_{1}+F_{2}+F_{12})-I(R+F_{1}+F_{2}+F_{12})),G_{mn})\\ =\varepsilon(\nabla(R-IR),\nabla G_{mn})+\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})\\ +\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})+\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})$ (4.34)
@@
Using Lemma (4.1 to 4.4), we have that when source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, then the diffusion term is bounded by
$(\nabla(u-Iu),\nabla G_{mn})_{\Omega}\leq C\frac{(\log(1/\varepsilon^{3/2})^{3}}{N}.$ (4.35)
Now, the second term of the R.H.S of (4.33) is $(b_{1}(Iu-u)_{x},G_{mn})$, using integration by parts, we get
$(b_{1}(Iu-u)_{x},G_{mn})=$ $(b_{1}(Iu-u),G_{mn})_{\partial\Omega}-(Iu-u,b_{1}G_{mn,x})_{\Omega}$
$(b_{1}(Iu-u)_{x},G_{mn})=$ $-(Iu-u,b_{1}G_{mn,x})_{\Omega}.$
and using Lemma (5.1 to 5.4), we have that when the source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$ then the convection term is bounded by
$(b_{1}(Iu-u)_{x},G_{mn})_{\Omega}\leq C\frac{\log^{2}(1/\varepsilon^{3/2})}{N}.$ (4.36)
Now, the third term of the R.H.S. of (4.33) is $(c(Iu-u),G_{mn})$ and using Lemma 5.5 we have when source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$ the reaction term bounds by
$(c(Iu-u),G_{mn})_{\Omega}\leq C\frac{\log^{2}(1/\varepsilon^{3/2})}{N}.$ (4.37)
Hence, combining (4.35), (4.36), and (4.37) completes the proof.
@@ 5 Numerical results and discussion
In this section, we take a numerical example to validate the theoretical results established in Theorems (4). We apply our method to a test problem to verify the uniform rate of convergence only in the $x$-layer and coarse layer regions.
@@ 5.1 Numerical implementation
$C^{N}_{\varepsilon}=\log_{2}\left(\frac{\|U_{N}-U_{2N}\|_{L^{\infty}(\Omega)}}{\|U_{2N}-U_{4N}\|_{L^{\infty}(\Omega)}}\right)$
@@
Example 5.1 We investigate the turning point singularly perturbed equation provided below
$\begin{split}-\varepsilon\Delta u-x(x^{2}+e^{1+xy})u_{x}+(3+x^{2}e^{x})u&=f(x,y),\quad\text{in}\;\Omega=(-1,1)\times(-1,1),\\ u&=0,\quad\Gamma=\partial\Omega\end{split}$ (5.1)
For this particular problem, the source term is set as $f(x,y)=\frac{xy}{1+x^{2}+y^{2}}$.
A key feature of this problem is its turning point at $x=0$, which gives rise to an interior layer along the $x$-axis. In parallel, two boundary layers are present at the bottom of the boundary y=1y=-1 and the top of the boundary $y=1$ of the domain. A double-mesh principle is used to improve the accuracy and convergence behavior of the numerical scheme. This scheme allows us to estimate errors and evaluate the rate of convergence for different values of the perturbation parameter $\varepsilon$ and the mesh discretization parameter N without knowing the exact solution. The resultant error values and computed rates of convergence are systematically presented in Tables.
@@ 5.2 Validation
For text example (5.1), Tables (1), (2), (4) and (5) display the numerical errors for subdomains $\Omega_{c}$, $\Omega_{f,x}$, $\Omega_{f,y}$ and $\Omega_{f,xy}$ respectively. Additionally, Tables (3) and (6) present the rates of convergence when the source point is located within $\Omega_{c}\cup\Omega_{f,x}$ and $\Omega_{f,y}\cup\Omega_{f,xy}$, respectively. As shown in Table (3), the proposed scheme is uniformly convergent for the source point that lies in $\Omega_{c}\cup\Omega_{f,x}$. Conversely, Table (6) indicates that uniformity is lost when the source point lies in $\Omega_{f,y}\cup\Omega_{f,xy}$. The solution plot clearly captures the interior and boundary layers. Furthermore, the numerical convergence rates, which are linear up to a logarithmic factor when the source point lies within the subdomain $\Omega_{c}\cup\Omega_{f,x}$, confirm our theoretical predictions.
@@ 6 Conclusions
In this article, we employ a standard finite element method to solve a two-dimensional singularly perturbed convection-diffusion turning point problem. Because the convection coefficient change sign within the domain, the problem has a interior layer along $x$-axis and two boundary layers along $y$-axis. Parameter uniform error estimates are obtained for the coarse and $x$ layer regions using Shishkin-type mesh with specific mesh transition parameter. Additionally, pointwise error estimates are established through the application of the discrete Green’s function. Stynes analysis indicated, uniformly convergent in the perturbation parameter of order $N^{-1}\log N$ in global energy norm and of order $N^{-1/2}\log^{3/2}N$ pointwise near the outflow boundary but not in the coarse region. We establish uniform convergence in the perturbation parameter of order $N^{-1}\log^{3}(1/\varepsilon^{3/2})$ pointwise in the coarse and $x$-layer regions.
@@ Acknowledgments
The authors are thankful to the anonymous referees for their valuable comments and suggestions to improve the paper. The authors thank LNMIIT, Jaipur for providing the facilities and support necessary to complete this research. The authors also acknowledge the DST-FIST program (Govt. of India) for providing the financial support for setting up the computing lab facility under the scheme ”Fund for Improvement of Science and Technology” (FIST - No. SR/EST/MS-1/2018/24). The authors also acknowledge the technical support provided by the LNM Institute of Information Technology, India.

### v2

@@ 5 Estimates related to convection and reaction terms
$\leq CN\|IF_{12}-F_{12}\|_{L^{\infty}(\Omega_{c})}\;\|G_{mn}\|_{\Omega_{c}}$
$\leq CN\sqrt{N\log\frac{1}{\varepsilon}}\left[\frac{\varepsilon^{-2}}{N^{2}}+\frac{\varepsilon^{-3/2}}{N^{2}}+\frac{1}{\varepsilon N^{2}}\right]\varepsilon^{\alpha+\frac{3\beta}{2}}$
$\leq C\varepsilon^{\alpha+\frac{3\beta}{2}-2}\frac{(\log(1/\varepsilon))^{1/2}}{N^{1/2}}\leq\frac{C(\log(1/\varepsilon))^{1/2}}{N^{1/2}}\quad\text{when}\;\alpha+\frac{3\beta}{2}\geq 2.$ (5.20)
Hence, when source point $(x_{m},y_{n})\in\Omega_{f,x}$, we have
$|(Iu-u,b_{1}G_{mn,x})|\leq C\frac{(\log(1/\varepsilon))^{1/2}}{N^{1/2}}.$
@@
(ii) Similarly when the source point $(x_{m},y_{n})\in\Omega_{c}$, using Lemma 3.1 yields $\|G_{mn}\|\leq CN$, we have
$|(Iu-u,b_{1}G_{mn,x})|\leq C.$
Remark 5.5.1
When the source point $(x_{m},y_{n})\in\Omega_{c}$, we can obtain only $\varepsilon$-independent bounds; the present approach does not provide a convergence rate as reported [23].
Following a similar argument to that discussed in the previous lemmas (5.2 - 5.5), we obtained the next Lemma.
Lemma 5.6
For $\alpha\geq 5/2$, $\beta\geq 2/3$, the following estimates holds:
  1. (i)
@@
    When the source point $(x_{m},y_{n})\in\Omega_{f,x}\;\text{then}\;|((c-b_{1,x}),(Iu-u)G_{mn})|\leq C\frac{\log(1/\varepsilon)}{N^{3/2}}$,
  2. (ii)
    when the source point $(x_{m},y_{n})\in\Omega_{c}\;\text{then}\;|((c-b_{1,x}),(Iu-u)G_{mn})|\leq C\frac{(\log(1/\varepsilon^{3/2}))^{7/2}}{N}.$
6 Main Result
Using the analysis from the preceding sections, we derive our main result, which confirms an $O(N^{-1/2}(\log(1/\varepsilon))^{1/2})$ rate of convergence within the $x$-layer region. For the coarse layer region, the result is $\varepsilon$-independent, but we could not determine a specific order of convergence.
Theorem 6.1
@@ 6 Main Result
Let $u_{h}$ denote the approximate solution of the problem 1.2 and the source point $(x_{m},y_{n})\in\Omega_{f,x}$, then for $\alpha\geq 5/2$ and $\beta\in[2/3,1]$ the following estimate hold
$|(u-u_{h})(x_{m},y_{n})|\leq C\frac{(\log(1/\varepsilon))^{1/2}}{N^{1/2}}.$
Proof. Let $(x_{m},y_{n})$ be a mesh point, then
$(Iu-u_{h})(x_{m},y_{n})=(\delta_{mn},Iu-u_{h}),$ (6.1)
@@
where $\delta_{mn}$ is the Kronecker delta function defined on mesh point $(x_{m},y_{n})$, then
$(Iu-u_{h})(x_{m},y_{n})$ $=(\delta_{mn},Iu-u_{h})=B(Iu-u_{h},G_{mn}),$
$\text{where}\quad B(u,v)$ $=\varepsilon\int_{\Omega}\nabla u.\nabla v\;dx+\int_{\Omega}b_{1}u_{x}v\;dx+\int_{\Omega}cuv\;dx$
$=\varepsilon(\nabla u,\nabla v)+(b_{1}u_{x},v)+(cu,v).$
$\implies(Iu-u_{h})(x_{m},y_{n})$ $=B(Iu-u_{h},G_{mn})$
$=B(Iu-u,G_{mn})+B(u-u_{h},G_{mn}).$
But Galerkin’s orthogonality implies $B(u-u_{h},G_{mn})=0,$ therefore
$(Iu-u_{h})(x_{m},y_{n})=B(Iu-u,G_{mn})$
$=\varepsilon\int_{\Omega}\nabla(Iu-u).\nabla G_{mn}\;dxdy+\int_{\Omega}b_{1}(Iu-u)_{x}G_{mn}\;dxdy+\int_{\Omega}c(Iu-u)G_{mn}\;dxdy$
$=\varepsilon(\nabla(Iu-u),\nabla G_{mn})+(b_{1}(Iu-u)_{x},G_{mn})+(c(Iu-u),G_{mn}).$
Applying integration by parts yields
$(Iu-u_{h})(x_{m},y_{n})=\varepsilon(\nabla(Iu-u),\nabla G_{mn})-(b_{1}(Iu-u),G_{mn,x})+((c-b_{1,x}),(Iu-u)G_{mn})$
therefore,
$|(Iu-u_{h})(x_{m},y_{n})|\;\leq|\varepsilon(\nabla(Iu-u),\nabla G_{mn})|+|(b_{1}(Iu-u),G_{mn,x})|+|((c-b_{1,x}),(Iu-u)G_{mn})|.$ (6.2)
To obtain the error estimate, we aim to estimate the term on the R.H.S. of (6.2). Consider the diffusion term of (6.2) and decompose the solution $u$ into regular and singular parts as described in Theorem 1.2. We obtain
$|\varepsilon(\nabla(u-Iu),\nabla G_{mn})|\;=|\varepsilon(\nabla((R+F_{1}+F_{2}+F_{12})-I(R+F_{1}+F_{2}+F_{12})),\nabla G_{mn})|\\ \leq|\varepsilon(\nabla(R-IR),\nabla G_{mn})|+|\varepsilon(\nabla(F_{1}-IF_{1}),\nabla G_{mn})|\\ +|\varepsilon(\nabla(F_{2}-IF_{2}),\nabla G_{mn})|+|\varepsilon(\nabla(F_{12}-IF_{12}),\nabla G_{mn})|$ (6.3)
@@
Using Lemmas (4.2 - 4.5), for the source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, then the diffusion term is bounded by
$|(\nabla(u-Iu),\nabla G_{mn})|\leq C\frac{(\log(1/\varepsilon^{3/2})^{7/2}}{N}.$ (6.4)
Consider the second term $|(b_{1}(Iu-u),G_{mn,x})|$ of (6.2). Applying Lemmas (5.2 - 5.5) for any source point $(x_{m},y_{n})\in\Omega_{f,x}$, we obtain
$|(Iu-u,b_{1}G_{mn,x})|\leq C\frac{(\log(1/\varepsilon))^{1/2}}{N^{1/2}}.$ (6.5)
Next consider $|((c-b_{1,x}),(Iu-u)G_{mn})|$, i.e. the third term of (6.2), Lemma 5.6 yields for any source point $(x_{m},y_{n})\in\Omega_{f,x}$, the following estimate hold
$|((c-b_{1,x}),(Iu-u)G_{mn})|\leq C\frac{\log(1/\varepsilon)}{N^{3/2}}.$ (6.6)
Combining (6.4), (6.5), and (6.6) yields the result.
Remark 6.1.1
Applying Lemmas (4.2 - 4.5), Lemmas (5.2 - 5.5), Lemma 5.6 and Green’s function estimate Lemma 3.1 for the source point $(x_{m},y_{n})\in\Omega_{c}$, we have $|(u-u_{h})(x_{m},y_{n})|\leq C.$ Here we obtain an $\varepsilon$-independent bound for the estimate; however no convergence is obtained and it agrees with the findings of [23].
Remark 6.1.2
For the coarse layer region, we have not yet found an order of convergence, but we believe higher-degree polynomials will help us to obtain it. We explore it further in future work.
@@ 7 Numerical results and discussion
In this section, we take a numerical example to validate the theoretical results established in Theorems (6.1). We apply our method to a test problem to verify the uniform rate of convergence only in the $x$-layer and coarse layer regions.
@@ 7.1 Numerical implementation
$C^{N}_{\varepsilon}=\log_{2}\left(\frac{\|U_{N}-U_{2N}\|_{L^{\infty}(\Omega)}}{\|U_{2N}-U_{4N}\|_{L^{\infty}(\Omega)}}\right).$
@@
Example 7.1
We investigate the turning point singularly perturbed equation provided below
$\begin{split}-\varepsilon\Delta u-x(x^{2}+e^{1+y^{2}})u_{x}+(2-x^{2}-y^{2})u&=f(x,y),\quad\text{in}\;\Omega=(-1,1)\times(-1,1)\\ u&=0,\quad\Gamma=\partial\Omega.\end{split}$
For this particular problem, the source term is set as $f(x,y)=\frac{xy}{1+x^{2}+y^{2}}$. A key feature of this problem is its turning point at $x=0$, which gives rise to an interior layer along the $x$-axis. In parallel, two boundary layers are present at the bottom of the boundary y=1y=-1 and the top of the boundary $y=1$ of the domain. A double-mesh principle is used to improve the accuracy and convergence behavior of the numerical scheme. This scheme allows us to estimate errors and evaluate the rate of convergence for different values of the perturbation parameter $\varepsilon$ and the mesh discretization parameter $N$ without requiring the exact solution. The resultant error values and computed rates of convergence are systematically presented in tables.
@@ 7.2 Validation
For text example (7.1), Tables (1), (2), (4) and (5) display the numerical errors for subdomains $\Omega_{c}$, $\Omega_{f,x}$, $\Omega_{f,y}$ and $\Omega_{f,xy}$ respectively. Additionally, Tables (3) and (6) present the rates of convergence when the source point is located within $\Omega_{c}\cup\Omega_{f,x}$ and $\Omega_{f,y}\cup\Omega_{f,xy}$, respectively. As shown in Table (3), the proposed scheme is uniformly convergent for the source point that lies in $\Omega_{f,x}$. Conversely, Table (6) indicates that uniformity is lost when the source point lies in $\Omega_{f,y}\cup\Omega_{f,xy}$. The solution plot clearly captures the interior and boundary layers. Furthermore, we obtain the numerical convergence rates, which are $N^{-1/2}$ up to a logarithmic factor when the source point lies within the subdomain $\Omega_{f,x}$, confirming our theoretical predictions.
@@ 8 Conclusions
In this article, we employ a standard finite element method to solve a two-dimensional singularly perturbed convection-diffusion problem with a turning point. Because the convection coefficient changes sign within the domain, the problem has an interior layer parallel to the $y$-axis at $x=0$ and two boundary layers parallel to the $x$-axis at y=1y=-1 and $y=1$. Parameter uniform error estimates are obtained for $x$-layer regions. For the coarse region, we obtain only $\varepsilon$-independent estimates, similar to those reported [23], which uses a Shishkin-type mesh with a specific mesh transition parameter. Additionally, pointwise error estimates are established through the application of the discrete Green’s function. We prove uniform convergence in the perturbation parameter of order $N^{-1/2}(\log(1/\varepsilon))^{1/2}$ pointwise in the $x$-layer region.
@@ Acknowledgments
The authors are thankful to the anonymous referees for their valuable comments and suggestions to improve the paper. The authors thank LNMIIT, Jaipur, for providing the facilities and support necessary to complete this research. The authors also acknowledge the DST-FIST program (Govt. of India) for providing the financial support for setting up the computing lab facility under the scheme ”Fund for Improvement of Science and Technology” (FIST - No. SR/EST/MS-1/2018/24). The authors also acknowledge the technical support provided by the LNM Institute of Information Technology, India.